OpenStudy (anonymous):

A cell phone plan offers a monthly rate of $25 for unlimited minutes plus $.05 per text message. Which algebraic expression below represents the total cost for this cell phone plan in terms of "x"?

OpenStudy (anonymous):

x=25+.05y where y is the number of test sent out

OpenStudy (anonymous):

smilz39 is right, but perhaps the Y is un-needed. The initial cost is 25 dollars Then you add the cost of the texts. This is the cost of each text multiplied by the number of texts sent. \[0.5 \times \chi\] Where x is equal to the number of calls Total costs = \[25 + 0.5\chi\] An example. George pays the 25 but also makes 30 texts \[ 25 + 0.5 \times 30 = 40\]